Complete the argument, giving a reason for each step.

AB= BC Given

AC= AB+BC A. ?

AC=AB+AB B. ?

AC=2(AB) C. ?

thank you

Answer by Edwin McCravy(20055)

Complete the argument, giving a reason for each step.
 
AB = BC   Given
 
AC = AB+BC  A. A whole is equal to the sum of its parts
 
AC = AB+AB  B. Equals may be substituted for each other and we
               are given that BC = AB
 
AC = 2(AB)  C. Because AB+AB=1*AB+1*AB= (1+1)*AB and 1+1=2   
 
Edwin